OpenStudy (anonymous):

Find the midpoint between points (20, 20) and (94, 40). Answer (57, 30) (30, 57) (20, 40) (40, 20)

OpenStudy (harsimran_hs4):

 get hold off how i got the mid point (x coordinate) from figure then we will see y

OpenStudy (anonymous):

Use the midpoint formula:\[\left( \frac{ x_{1}+x_{2} }{ 2 }, \frac{ y_{1}+y_{2} }{ 2 }\right)\]

OpenStudy (harsimran_hs4):

then if you apply not geometry u will find ymid = (y + b)/2 so in general when you need to find mid point of two points just take average i.e (1st + 2nd )/2 for both x and y

OpenStudy (harsimran_hs4):

in your ques x = (94+20)/2 = 57 y = (20+40)/2 = 30 (57, 30)